# Prompt to Pwn: Automated Exploit Generation for Smart Contracts

### v1

@@ Abstract
We explore the feasibility of using LLMs for Automated Exploit Generation (AEG) against vulnerable smart contracts. We present ReX, a framework integrating LLM-based exploit synthesis with the Foundry testing suite, enabling the automated generation and validation of proof-of-concept (PoC) exploits. We evaluate five state-of-the-art LLMs (GPT-4.1, Gemini 2.5 Pro, Claude Opus 4, DeepSeek, and Qwen3 Plus) on both synthetic benchmarks and real-world smart contracts affected by known high-impact exploits. Our results show that modern LLMs can reliably generate functional PoC exploits for diverse vulnerability types, with success rates reaching up to 92%. Notably, Gemini 2.5 Pro and GPT-4.1 consistently outperform others in both synthetic and real-world scenarios. We further analyze factors influencing AEG effectiveness, including model capabilities, contract structure, and vulnerability types. We also collect the first curated dataset of real-world PoC exploits to support future research.
Introduction
Once smart contracts were deployed, even a small vulnerability may be permanently exploitable, resulting in substantial financial losses. We show an example.
In February 2025, a smart contract exploit on Bybit’s Safe multi-signature wallet allowed attackers to upgrade the contract implementation and drain 1.5 billion US dollars (Rommen and Sehmbi 2025).
To address the security issues, many vulnerability detection tools have been proposed, including Slither (Feist, Grieco, and Groce 2019), Mythril (Nikolić et al. 2018), and Oyente (Luu et al. 2016), but existing static and symbolic analysis tools suffer from low accuracy and limited scalability and perform poorly in practice (Sendner et al. 2024).
Large language models (LLMs) have demonstrated impressive capabilities across a wide range of applications, including code-related tasks such as generation (Laurie et al. 2022), summarization, and bug fixing (Pearce et al. 2022). Not surprisingly, LLMs also hold strong potential for assisting in the identification of smart contract vulnerabilities (Chen et al. 2025; David et al. 2023; Xiao et al. 2025).
We step further to explore their effectiveness in generating verifiable exploits. We focus on a few research questions:
  • RQ1: How well can LLMs perform automated exploit generation (AEG) for smart contract vulnerabilities on both benchmarks and real-world attack data?
@@ Introduction
To investigate whether LLMs can move beyond detection to AEG and validate executable proof-of-concept (PoC) exploits for smart contracts, we propose ReX, a new framework that integrates LLM-based exploit synthesis with the Foundry testing environment, enabling end-to-end exploit generation, compilation, execution, and verification. We evaluate five most latest LLMs (GPT-4.1, Gemini 2.5 Pro, Claude Opus 4, DeepSeek-R1, Qwen3 Plus) on their ability to generate full exploit contracts and test scripts, targeting eight distinct real-world vulnerability types (e.g., reentrancy, integer overflows, improper access control).
To further improve output automation, we design three optimization techniques, including prompt refactoring, compiler feedback loops, and templated test harness generation. These enhancements significantly improve success rates for underperforming models. Our results show that:
  • GPT-4.1, Gemini 2.5 Pro, and Claude Opus 4 achieve success rates 80%+ on most vulnerability categories, with GPT-4.1 peaking at 92.5% on arithmetic bugs.
  • DeepSeek-R1 and Qwen3 Plus produce valid exploits in 60–70% of cases but show inconsistency across more complex vulnerability types.
  • Over 76% of generated artifacts pass compilation and test execution end-to-end without human intervention.
In addition, we curate the first dataset111provided upon request. of manually written PoC exploits targeting real-world contracts, consisting of 38+ audited attack cases from expert security researchers.
Finally, we validate LLM-generated exploits on smart contracts that suffered from historical high-impact attacks. In 4 out of 10 cases, Gemini 2.5 Pro was able to independently rediscover valid exploits. It not only demonstrated autonomous exploit discovery capabilities, but also exhibited expert-like reasoning strategies similar to those employed by seasoned smart contract auditors.
To facilitate reproducible research and open science, we provide the code, including the detailed prompts and raw experiment data as supplementary materials, and will open-source them upon paper acceptance.
Technical Warmups
Smart contract vulnerabilities.
@@
Smart contracts are vulnerable to bugs and exploits (Tolmach et al. 2021). Once deployed, contracts cannot be altered, leaving any flaws permanently exposed. To standardize and detect such issues, the SWC Registry (swc 2018) categorizes common vulnerabilities like reentrancy (SWC-107), integer overflows (SWC-101), and access control flaws (SWC-105), forming the basis for many tools and benchmarks. As decentralized applications grow in complexity, attackers increasingly leverage sophisticated techniques such as arbitrage (Torres, Steichen et al. 2019; Cernera et al. 2023), transaction order manipulation (Eskandari, Moosavi, and Clark 2019), and proxy-based upgrade attacks (Meisami and Bodell III 2023).
LLMs.
@@
We evaluate five latest LLMs in this work. GPT-4.1 (OpenAI) is known for strong reasoning and consistent performance across code and language tasks. Gemini 2.5 Pro (Google) handles text, code, and images well, matching GPT-4-class performance. Claude Opus 4 (Anthropic) offers long-context reasoning and is suited for safety-critical tasks. DeepSeek-V2 delivers high performance in bilingual and code tasks. Qwen3 Plus (Alibaba) provides strong code and dialogue capabilities, especially in English and Chinese.
Contract testing suite.
@@
We use Foundry, a Rust-based toolchain for smart contract development and testing (Fou 2025). Foundry supports dependency management, compilation, deployment, and testing. Its built-in test runner (forge test) enables efficient, large-scale validation of Solidity exploits, outperforming tools like Hardhat and Truffle.
Static analytic tools.
@@
We use two. SUMO is a static analysis tool for smart contracts that estimates code understandability through a composite complexity score (Barboni, Morichetta, and Polini 2022). It captures Solidity-specific features like control-flow depth, nesting, and branching. Slither is another widely used static analysis tool for detecting vulnerabilities and computing code metrics in Solidity contracts (Feist, Grieco, and Groce 2019). It provides accurate and fine-grained structural insights.
ReX and Datasets
Figure 1: Overview of the ReX framework
As shown in Figure 1, ReX is a fully automated pipeline and proceeds in five steps:
@@
Step 2: script generation. Gven a vulnerable smart contract, through carefully designed prompt, the LLM will generate two related Foundry scripts: the first is an exploit contract designed to exercise the vulnerable code path; while the second is a test contract that validates whether the exploit succeeds under some specific conditions.
We enabled the LLMs to iteratively optimize its prompts (Chen et al. 2025; Grubisic et al. 2024; Sepidband et al. 2025). As a result, it is capable of generating exploit and test scripts with intact import paths, designed to be as compilable and executable as possible. At the same time, we enabled the model’s reasoning mode by guiding it to reason step by step, aiming to improve the accuracy of its responses.
@@
Step 5: Iterative feedback loop If compilation or testing fails, the error messages and scripts are returned to the LLM, which attempts to correct and regenerate them. This loop repeats until a valid exploit is found or a maximum retry limit is reached. This feedback-driven refinement is inspired by prior success in LLM-based program synthesis (Bi et al. 2024; Grubisic et al. 2024), and significantly boosts the success rate of exploit generation.
What datasets used for evaluation? We employ two datasets to evaluate ReX’s ability: one well-established benchmark for controlled testing (Durieux et al. 2020), and one curated collection of real-world attack cases.
@@
  • As a classical smart contract vulnerability dataset. It has been widely used in both smart contract vulnerability detection community and related researches (Ferreira et al. 2020)(Chen et al. 2025)(Durieux et al. 2020);
@@
  • The simplicity and clarity of the dataset make it particularly suitable for the initial evaluation of LLM’s capability to generate vulnerability exploitation PoCs. Smartbugs-curated provides a collection of vulnerable Solidity smart contracts organized according to the DASP taxonomy (Chen et al. 2025).
@@
Evaluating AEGs
Exp1: Benchmarking on SmartBugs-Curated.
The first experiment (Exp1) provides a baseline assessment to evaluate LLM’s performance in AEG with the structurally simple SmartBugs-Curated dataset.
@@
We manually check the consistency of each smart contract before and after version migration, and exclude contracts that do not meet the criteria. In fact, we found that the vast majority of smart contracts only required updating the compiler version. With version migration, these contracts also differ from their original forms, which helps reduce the likelihood that they appear in the LLM’s training data.
Next, we use the previously mentioned evaluation framework to generate and assess AEG.
For specific vulnerability types, we do not only rely on Foundry test results for evaluation. LLMs may generate exploits that trigger overflows or unexpected reverts-especially in denial-of-service and reentrancy cases—causing the tests to fail. Despite of it, these cases still reflect successful and valid exploit attempts.
Table 1: AEG Performance for Smart Contract Vulnerabilities against of five mainstream LLMs
@@
Specifically, Gemini 2.5 Pro achieved the highest average success rate (67.3%) across diverse vulnerability types, excelling in arithmetic (92.9%), front running (75.0%), and unchecked low-level calls (56.7%). GPT-4.1 followed with 58.1%, showing consistent results in arithmetic (85.7%) and time manipulation (80.0%). Claude Opus 4 ranked third (63.3%) but showed strong robustness across all categories, including DoS (100.0%) and access control (85.6%). In contrast, DeepSeek and Qwen3-plus exhibited noticeably lower success rates (48.3% and 28.8%, respectively), struggling especially with complex vulnerabilities such as reentrancy and unchecked low-level calls.
Exp2: Real-world exploits with Web3-AEG.
@@
The contract uses block data as its randomness source, making outcomes predictable. Attackers win every bet by predicting correct numbers off-chain (SlowMist1 2024).
@@
The contract fails to verify the caller in its handler onERC1155Received, allowing arbitrary minting of ERC20 tokens. This vulnerability resulted in a cumulative losses exceeding 138.78 ETH (SlowMist2 2024).
@@
The Pine Protocol uses a shared vault for both legacy and upgraded lending pool contracts. This leads to flashloan-based reentrancy exploit (Mutual 2024).
@@
Answer to RQ1: can LLMs perform AEGs?
@@
Exploring Key Factors Influencing LLM Performance in AEG
We now turn to investigate what determines their success or failure in AEG. Despite strong performance overall, the effectiveness of LLMs is inconsistent across contracts. Some highly complex contracts are easily exploited, while others with simpler structure resist attack. We discuss four factors.
Factor 1: LLM capabilities and failure patterns.
@@
Table 2: Five LLMs on Code Generation Benchmarks
@@
[1pt/1pt] Qwen3 61.8% 1291 54.2%
As shown in Table 2, Gemini 2.5 Pro outperforms other models on two widely used programming benchmarks (RankedAGI 2024). These scores correlate with its superior AEG performance in our experiments, reinforcing the view that general coding ability is predictive of LLMs’ exploit generation capacity. Similarly, Claude Opus 4, which also achieved high scores across standard coding benchmarks, demonstrated strong performance in our AEG experiments as well.
Moreover, we identified two recurring failure patterns (P).
  • P1: cryptographic limitations. Many LLMs incorrectly generate non-checksummed Ethereum addresses due to training data biases. LLMs cannot compute Keccak-256 hashes needed for EIP-55 compliance.
  • P2: semantic misunderstandin. LLMs frequently mishandle the ‘payable‘ modifier, omitting necessary casts or using it inconsistently. This reflects a deeper challenge in enforcing compiler-level semantic constraints.
Both success and failure reflect the model’s internal understanding. LLMs function as probabilistic pattern matchers, not reasoning engines. Their systematic errors highlight the boundaries of what current LLMs can achieve in AEG.
Factor 2: properties of target contracts
We next assess whether structural properties of the contracts influence AEG outcomes.
Using the Web3-AEG dataset, we extract source-level metrics such as nSLOC, complexity score, and external call count (Figure 4) with analyses via Cramér's V (Table 4). We also listed main factors of exploited contracts in Table 3.

### v2

@@ Abstract
Smart contracts are important for digital finance, yet they are hard to patch once deployed. Prior work mostly studies LLMs for vulnerability detection, leaving their automated exploit generation (AEG) capability unclear. This paper closes that gap with ReX, a framework that links LLM-based exploit synthesis to the Foundry stack for end-to-end generation, compilation, execution, and verification. Five recent LLMs are evaluated across eight common vulnerability classes, supported by a curated dataset of 38+ real incident PoCs and three automation aids: prompt refactoring, a compiler feedback loop, and templated test harnesses. Results indicate strong performance on single-contract PoCs and weak performance on cross-contract attacks; outcomes depend mainly on the model and bug type, with code structure and prompt tuning contributing little. The study also surfaces gaps in current defenses against LLM-driven AEG, pointing to the need for stronger protections.
1 Introduction
Smart contracts are widely used in digital finance. Their immutability is a double-edged sword: it builds trust in stored data, but once deployed, even small flaws can be exploited indefinitely, causing large losses. For example, in February 2025, an exploit of Bybit’s Safe multi-signature wallet let attackers upgrade the contract and drain about US\$1.5 billion [1].
To improve smart contract security, many detection tools have been developed, including Slither [2], Mythril [3], and Oyente [4]. However, static and symbolic analyses often have low accuracy, limited scalability, and weak real-world performance [5]. Recent large language models (LLMs) show strong capabilities on code tasks, such as generation [6], summarization, and bug fixing [7]. Thus, leveraging the code-processing capabilities of LLMs is a promising direction for detecting smart contract vulnerabilities [8, 9, 10].
While LLMs have been studied for detecting smart contract vulnerabilities, their ability to generate exploits is far less explored. This gap matters for two reasons. First, exploit generation can confirm a vulnerability and gauge its severity. Second, because LLMs are easy to access via APIs, we must assess how readily novice attackers (“script kiddies”) could use them to launch real attacks and harm the smart contract ecosystem.
To fill the research gap, we propose the first empirical study to understand the automated exploit generation (AEG) capabilities of LLMs. In this study, we focus on three research questions:
  • $\bullet$
    RQ1: How well can LLMs perform AEG for smart contract vulnerabilities on benchmarks and real-world attack data?
  • $\bullet$
@@ 1 Introduction
To conduct the empirical study and answer the research questions, we present ReX, a framework that couples LLM-based exploit synthesis with the Foundry testing stack for end-to-end generation, compilation, execution, and verification. We assess five recent LLMs (GPT-4.1, Gemini 2.5 Pro, Claude Opus 4, DeepSeek-R1, Qwen3 Plus) on producing complete exploit contracts and test scripts across eight real-world vulnerability classes (e.g., reentrancy, integer overflow, improper access control). To boost automation, we add three techniques—prompt refactoring, a compiler-in-the-loop feedback cycle, and templated test-harness generation—which markedly raise success rates for weaker models. Beyond the framework, we curate the first dataset [^1] To support reproducibility, we include code, prompts, and raw data as supplementary materials and will open-source them upon acceptance. of hand-written PoC exploits for real contracts [^2] Ethics: All contracts in Web3-AEG come from public incident reports (e.g., SlowMist, CertiK), with original PoCs already disclosed. No new undisclosed exploits are released., covering 38+ expert-audited attack cases. Finally, we validate LLM-generated exploits on contracts involved in past high-impact incidents.
Our study shows that LLMs can generate working proof-of-concept (PoC) exploits for single-contract bugs with high success and good speed, but they struggle with cross-contract cases. The effectiveness of AEG depends mainly on the underlying LLM and the bug type; code structure and prompt tweaks have limited impact. Finally, we reveal gaps in current defenses against LLM-based AEG, indicating a need for stronger protections.
In summary, our contributions are:
  • $\bullet$
    To our knowledge, this is the first systematic empirical study of LLMs’ automated exploit generation (AEG) capabilities.
  • $\bullet$
    We report practical strengths and limits of LLMs for AEG, identify key factors that affect performance, and derive implications for defense.
  • $\bullet$
    We implement the ReX framework to evaluate LLM-based smart-contract AEG and curate a public dataset [^3] Git repo: https://github.com/Crimson798/AEG-PoC of exploits for both synthetic and real-world vulnerabilities.
2 Background
@@
Smart contract vulnerabilities. Smart contracts are vulnerable to bugs and exploits [11]. Once deployed, contracts cannot be altered, leaving any flaws permanently exposed. To standardize and detect such issues, the SWC Registry [12] categorizes common vulnerabilities like reentrancy (SWC-107), integer overflows (SWC-101), and access control flaws (SWC-105), forming the basis for many tools and benchmarks. As decentralized applications grow in complexity, attackers increasingly leverage sophisticated techniques such as arbitrage [13, 14], transaction order manipulation [15], and proxy-based upgrade attacks [16].
@@
LLMs. We introduce five LLMs. GPT-4.1 (OpenAI) is known for strong reasoning and consistent performance across code and language tasks. Gemini 2.5 Pro (Google) handles text, code, and images well, matching GPT-4-class performance. Claude Opus 4 (Anthropic) offers long-context reasoning and is suited for safety-critical tasks. DeepSeek-R1 delivers high performance in bilingual and code tasks. Qwen3-Plus (Alibaba) provides strong code and dialogue capabilities, especially in English and Chinese.
@@
Contract testing suite. Foundry is a Rust-based toolchain for smart contract development and testing [17]. Foundry supports dependency management, compilation, deployment, and testing. Its built-in test runner (forge test) enables efficient validation of Solidity exploits, outperforming tools like Hardhat/Truffle.
@@
Static analytic tools. We introduce two representative tools. SUMO is a static analysis tool for smart contracts that estimates code understandability through a composite complexity score [18]. It captures Solidity-specific features like control-flow depth, nesting, and branching. Slither is another widely used static analysis tool for detecting vulnerabilities and computing code metrics in Solidity contracts [2]. It provides accurate and fine-grained structural insights.
VulnerableContractPreprocessing(Step 1)LLM Prompting(Exploit + Test, Step 2)Auto Fixes(Step 3)Foundry Build & Test(Step 4)Failure?LLM Feedback Loop(Step 5)Input CleaningCode SynthesisScript NormalizationDeterministic Validation
Figure 1: End-to-end workflow of ReX. The system iteratively synthesizes, repairs, and validates exploit PoCs until Foundry confirms a working attack.
3 Approach
ReX is an end-to-end, automated pipeline for transforming a vulnerable Solidity contract into a validated proof-of-concept exploit. Our guiding principles are (i) automation to minimize human intervention across generation, compilation and validation stages; (ii) robustness to tolerate minor model inaccuracies through post-processing and feedback; and (iii) reproducibility to integrate with a deterministic testing stack (i.e., Foundry) to ensure consistent verification.
We adopt Foundry rather than Hardhat/Truffle for three reasons. First, Foundry provides deterministic replay and trace-level debugging, making PoC verification reproducible across runs. Second, its lightweight compilation pipeline greatly reduces end-to-end AEG latency. Third, cheat codes (e.g., vm.warp, vm.expectRevert) allow precise instrumentation of safety invariants without modifying the target contract. It is essential for contract-agnostic validation.
3.1 Design of ReX
ReX consists of a sequence of modular components (Figure 1) that perform contract preprocessing, LLM-driven exploit and test synthesis, optional script repairs, automated compilation/testing, and iterative refinement.
Below we describe these five steps in detail.
@@ 3.1 Design of ReX
Step 2: script generation. Given a vulnerable smart contract and a carefully designed prompt, the LLM will generate two related Foundry scripts: the first is an exploit contract designed to exercise the vulnerable code path; while the second is a test contract that validates whether the exploit succeeds under some specific conditions.
We enabled the LLMs to iteratively optimize their prompts [8, 19, 20]. As a result, it is capable of generating exploit and test scripts with intact import paths, designed to be as compilable and executable as possible. At the same time, we enabled the model’s reasoning mode by guiding it to reason step by step, aiming to improve the accuracy of its responses.
The full prompt set, e.g., base system prompt, exploit-generation prompt, test-generation prompt, and error-guided feedback loop, is in Appendix 0.A.
@@
Step 5: Iterative feedback loop. If compilation or testing fails, the error messages and scripts are returned to the LLM, which attempts to correct and regenerate them. This loop repeats until a valid exploit is found or a maximum retry limit is reached. This feedback-driven refinement is inspired by prior success in LLM-based program synthesis [21, 19], and significantly boosts the success rate of exploit generation.
3.2 Datasets for Evaluation
We employ two datasets to evaluate ReX: one well-established benchmark for controlled testing [22], and one curated collection of real-world attack cases.
@@ 3.2 Datasets for Evaluation
  • $\bullet$
    As a classical smart contract vulnerability dataset, SmartBugs-Curated has been widely used in both smart contract vulnerability detection community and related research [23][8][22];
  • $\bullet$
@@
  • $\bullet$
    The simplicity and clarity of the dataset make it particularly suitable for the initial evaluation of LLM’s capability to generate vulnerability exploitation PoCs. Smartbugs-curated provides a collection of vulnerable Solidity smart contracts organized according to the DASP taxonomy [8].
@@
We also stress its representativeness. The 38 real-world incidents in Web3-AEG span five years, seven chains, and cover all major vulnerability categories reported by SlowMist, CertiK, Immunefi, and public post-mortems. The median contract size (225 SLOC) and prevalence of multi-call interactions reflect typical DeFi deployments rather than synthetic toy examples. This diversity allows our evaluation to approximate real attacker conditions more closely.
Additional dataset statistics and construction details for both SmartBugs Curated and Web3-AEG are given in Appendix 0.C.
3.3 Threat Model
Our attacker is a non-privileged party equipped with local execution resources and public RPC access. The adversary is allowed to (i) fork historical on-chain state for off-chain reasoning, (ii) repeatedly query an LLM and run Foundry-based tests, and (iii) deploy arbitrary auxiliary contracts. The attacker does not possess private keys, oracle manipulation ability, MEV privileges, or access to privileged off-chain infrastructure.
An exploit is considered successful if the LLM-generated PoC deterministically violates at least one class-level safety invariant under a clean Foundry sandbox. This definition aligns with practical exploitability while avoiding over-claiming economic losses.
3.4 Safety Invariants for Automated PoC Validation
To avoid any contract-specific manual intervention, we define class-level safety invariants for each vulnerability category rather than per-contract invariants. These invariants are automatically checked by the Foundry test harness:
  • $\bullet$
    Unauthorized state change: balance increase, mint, or withdraw without satisfying the intended access control (reentrancy, access control).
  • $\bullet$
    Arithmetic correctness: violation of arithmetic constraints inside unchecked{} blocks that should not overflow or underflow (arithmetic).
  • $\bullet$
    Predictability success: ability to consistently compute outcomes above a fixed threshold (bad randomness).
  • $\bullet$
    Unexpected revert / state freeze: triggering persistent inability to progress the contract state (denial-of-service).
  • $\bullet$
    Temporal invariant break: manipulating timestamp or block.* fields to violate a time-based constraint (time manipulation).
These invariants are generic, contract-agnostic, and require no prior knowledge of the real exploit logic.
Here, we note that a key design choice is the use of class-level contract-agnostic safety invariants, instead of manual per-contract rules. We argue this is sufficient because (i) all eight vulnerability classes exhibit invariant-breaking symptoms observable from the victim’s external interface (balance changes, reverts, predictability, or temporal guards), and (ii) Foundry ensures that these symptoms are captured at the VM boundary exactly once per test run. Across 56 benchmark contracts and 38 real-world cases, we found no instance where a successful exploit failed to violate at least one class-level invariant.
Complete class-level safety invariants and their Foundry assertion encodings are included in Appendix 0.B.
4 Evaluation
4.1 Exp1: Benchmarking on SmartBugs-Curated.
The first experiment (Exp1) provides a baseline assessment to evaluate LLMs’ performance in AEG with the structurally simple SmartBugs-Curated dataset.
@@ 4 Evaluation
We manually check the consistency of each smart contract before and after version migration, and exclude contracts that do not meet the criteria. We found that the vast majority of smart contracts only required updating the compiler version. With version migration, these contracts also differ from original forms, which helps reduce the likelihood that they appear in the LLM’s training data.
Next, we use the previously mentioned evaluation framework to generate and assess AEG. For specific vulnerability types, we do not rely solely on Foundry test results for evaluation. LLMs may generate exploits that trigger overflows or unexpected reverts, especially in denial-of-service and reentrancy cases, causing the tests to fail. Despite this, these cases still reflect valid exploit attempts.
Table 1: AEG Performance for Smart Contract Vulnerabilities
@@
Gemini 2.5 Pro achieved the highest average success rate (67.3%) across diverse vulnerability types, excelling in arithmetic (92.9%), front running (75.0%), and unchecked low-level calls (56.7%). GPT-4.1 followed with 58.1%, showing consistent results in arithmetic (85.7%) and time manipulation (80.0%). Claude Opus 4 ranked third (63.3%) but showed strong robustness across all categories, including DoS (100.0%) and access control (55.6%). In contrast, DeepSeek and Qwen3-plus exhibited lower success rates (48.3% and 28.8%), struggling with complex vulnerabilities such as reentrancy and unchecked low-level calls.
4.2 Exp2: Real-world exploits with Web3-AEG.
@@
The contract uses block data as its randomness source, making outcomes predictable. Attackers win every bet by predicting correct numbers off-chain [24].
  • $\bullet$
@@
The contract fails to verify the caller in its handler onERC1155Received, allowing arbitrary minting of ERC20 tokens. This vulnerability resulted in a cumulative losses exceeding 138.78 ETH [25].
  • $\bullet$
@@
The Pine Protocol uses a shared vault for both legacy and upgraded lending pool contracts. This leads to flashloan-based reentrancy exploit [26].
  • $\bullet$
@@
4.3 Answer to RQ1: can LLMs perform AEGs?
@@
5 Key Factors Affecting LLM Performance
We now examine which factors determine success or failure in AEG. Although modern LLMs achieve strong results overall, their performance varies substantially across contracts. Some contracts with complex control flow are exploited easily, whereas other contracts with simpler logic resist attack. To understand this variation, we evaluate four dimensions: (1) intrinsic model capability, (2) target contract structure, (3) vulnerability type, and (4) prompt design.
To ground the discussion, we formalize what constitutes AEG success.
Definition 1(AEG success).
Let $C$ denote a target contract and let $E$ denote an LLM-generated exploit and test pair. A sample is counted as a successful AEG instance if the generated code compiles under Foundry and deterministically violates at least one class-level safety invariant described in Appendix 0.B.
This definition captures semantic exploitability in a controlled setting without requiring economic loss or network-level manipulation.
5.1 Factor 1: LLM capabilities and failure patterns
@@
Table 2: Code Generation Benchm.
@@
[1pt/1pt] Qwen3-plus 61.8% 1291 54.2%
Table 3: Impact of Iterative Repair
Vulnerability 1 turn Max.4 turns
Reentrancy 33% 71%
[1pt/1pt] Access Control 28% 65%
[1pt/1pt] Arithmetic 42% 59%
[1pt/1pt] DoS 18% 44%
[1pt/1pt] Time Manipulation 12% 39%
[1pt/1pt] Overall 27% 58%
As shown in Table 3, Gemini 2.5 Pro outperforms the other models on two widely used programming benchmarks [27]. These scores correlate with its superior AEG performance in our experiments, which supports the view that general coding ability is predictive of an LLM’s exploit generation capacity. Claude Opus 4, which also achieves high scores on standard coding benchmarks, likewise shows strong AEG performance.
We further examine how models fail. We identify two recurring failure patterns (P) that appear across datasets.
  • $\bullet$
    P1: cryptographic limitations. Many LLMs incorrectly generate non checksummed Ethereum addresses and cannot compute Keccak–256 hashes needed for EIP–55 compliance. These mistakes arise from token level prediction limits rather than any particular contract.
  • $\bullet$
    P2: semantic misunderstanding. LLMs frequently mishandle the payable modifier, omit necessary casts, or confuse state changing functions with view functions. This reflects a deeper challenge in enforcing compiler level semantic constraints for EVM programs.
Across all failed attempts, we classify root causes into four categories: syntactic Solidity errors (42%), incorrect or incomplete Foundry test harnesses (33%), unmet safety invariants (15%), and mistakes in reasoning about contract state and control flow (10%). This error distribution suggests that AEG failure is mainly driven by the model’s internal reasoning limits, rather than by contract complexity alone. We also compare single turn and multi turn generation. ReX can invoke an iterative repair loop where compilation errors or failing assertions are fed back to the model.
Table 3 shows that iterative repair almost doubles the overall success rate. In many cases the first attempt already contains a plausible exploit plan but misses one or two syntactic or semantic details. The feedback loop helps the model correct these mistakes and assemble a complete exploit and test pair.
Both success and failure reflect the model’s internal understanding. LLMs behave as probabilistic pattern matchers rather than fully reliable reasoning engines. Their systematic errors highlight the boundaries of what current LLMs can achieve in AEG.
5.2 Factor 2: properties of target contracts
We next assess whether structural contract properties influence AEG outcomes.
Using the Web3–AEG dataset, we extract source level metrics such as nSLOC, complexity score, and external call count, and we analyze their associations with AEG success via Cramér's V (Table 5). We also list representative complex contracts in Table 5.